# Bidirectional Nonlinear Optical Tomography: Unbiased Characterization of Off- and On-Chip Coupling Efficiencies

### v1

@@ II Outlook
In summary, BNOT provides a platform-agnostic approach to calibrate asymmetric coupling efficiencies in nonlinear PICs. BNOT works with SHG and squeezed-light measurements, so it can be deployed on existing setups without extra hardware to benchmark chip-to-fiber and inter-stage interfaces on the chip. An exciting near-term application of BNOT can be in high-device-yield nonlinear PICs experiments, such as recent demonstrations of wafer-scale on-chip multi-harmonic generation [15] and on-chip squeezing [12]. Beyond our exemplary investigation of on-chip squeezing and SHG, BNOT provides an unbiased calibration approach for other on-chip nonlinear phenomena [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 19, 20, 21, 22, 23, 3, 16].
@@
Looking ahead, an intriguing direction may be harnessing BNOT in adaptive control and self-calibrating architectures. By combining real-time forward/backward measurements with Bayesian [47, 48, 49, 50] or machine-learning estimators [51], one can automate coupling-loss compensation across large photonic arrays and dynamically track degradation or drift. Such closed-loop implementations may convert coupling characterization from a post-hoc diagnostic into an active subsystem of photonic hardware, supporting reproducible, system-level benchmarks as nonlinear PICs transition from laboratory demonstrations to scalable quantum and classical technologies.
III Acknowledgments
@@
The authors thank Shi-Yuan Ma, Mengjie Yu, Avik Dutt, and Sri Krishna Vadlamani for fruitful discussions. This work was supported by funding from the DARPA INSPIRED program.
IV Appendix
MC simulation parameters $d_{\text{eff}}$ $19.5$ pm/V Thin-film LiN effective nonlinear coefficient $L$ 66 mm  Length of PPLN waveguide AA $1.4$ µm2 Cross-sectional area of PPLN waveguide $\omega$ $193$ THz Pumping frequency (or $2\omega$)  (or $386$ THz)   at $1550$ nm (or $775$ nm) $n_{\omega}$ $2.18$  Refractive index of LiN (or $n_{2\omega}$) $2.14$  at $\omega$ (or $2\omega$) $\sigma_{\text{sqz}}$ $0.1\penalty 10000\ \text{dB}$  Squeezing (or SHG) (or $\sigma_{\text{shg}}$)    (or $20\penalty 10000\ \text{\%/W/cm${}^{2}$}$)   measurement noise STD $P_{2\omega}$ $100$ mW   Off-chip mean pump (or $P_{\omega}$) (or $10$ mW) power in squeezing (or SHG) experiment Off-chip pump power $\delta_{1}$ $0.03$ half of the fluctuation (or $\delta_{2}$) (or $0.05$) rate in squeezing (or SHG) experiment $\Delta k$ 0 m-1 Wave-vector mismatch
Table 1: MC simulation parameters of PPLN waveguide in squeezing and SHG experiments. STD: standard deviation.
IV.1 Efficiency thresholds for off-chip squeezing
@@
On-chip squeezing levels approaching $15\penalty 10000\ \text{dB}$ are particularly relevant for applications such as CV quantum error correction and precision quantum metrology [26, 39]. As shown in Fig. 5(a), achieving this on-chip benchmark requires an input coupling efficiency $\eta^{(2\omega)}_{1}\gtrsim 0.65$, which is attainable using current photonic integration technologies. For example, edge-coupled LiN modulators have demonstrated per-interface coupling efficiencies below $0.5\penalty 10000\ \text{dB}$ (corresponding to efficiencies above $90\penalty 10000\ \%$) at $1550\penalty 10000\ \text{nm}$ [52], indicating that such coupling efficiencies are experimentally feasible.

### v2

@@ IV Conclusion and outlook
In summary, BNOT provides a platform-agnostic method to calibrate asymmetric coupling efficiencies in nonlinear PICs. As we prototypically demonstrated for SHG and squeezed-light measurements, BNOT integrates seamlessly with existing experimental setups and enables accurate benchmarking of both chip-to-fiber interfaces and inter-stage connections. This capability is particularly valuable for high–device-yield nonlinear PIC platforms, including recent wafer-scale demonstrations of on-chip multi-harmonic generation [15] and on-chip squeezing [12], where device statistics and variations make unbiased coupling calibration essential. Beyond the two processes highlighted in this work, the same strategy can be applied to a broad class of nonlinear interactions such as SPDC [3, 4, 5], FWM [6, 7, 8, 9, 10, 11, 12, 13], OPO [19, 20, 21, 22] and OPA [23, 3, 16], providing a unified approach for coupling-resolved benchmarking as integrated nonlinear photonics continues to scale.
Conceptually, the complementary SHG and SPDC mechanisms used here reflect a broader principle: nonlinear processes introduce measurement asymmetries that break the intrinsic $\eta_{1}\eta_{2}$ degeneracy that limits purely linear transmission-based calibration methods. In practice, any bidirectional or multi-process approach naturally exhibits small forward and backward differences, for example, in phase matching or mode overlap. These effects can similarly be utilized for BNOT, given that the essential requirement is a controlled asymmetry in the forward model that distinguishes the two interfaces in a physically meaning way. When this asymmetry is taken into account, the joint efficiency inference converges reliably to the true coupling values.
@@
Looking forward, a promising direction is to integrate BNOT-style tomography into adaptive control and self-calibrating photonic architectures. Real-time forward and backward measurements, combined with Bayesian [47, 48, 49, 50] or machine-learning estimators [51], may enable automatic compensation of coupling losses across large photonic arrays and continuous tracking of drift or degradation. Such closed-loop implementations have the potential to transform coupling calibration from a post-measurement diagnostic into an active subsystem of photonic hardware, supporting reproducible, system-level benchmarks as nonlinear PICs transition from laboratory demonstrations to scalable quantum and classical technologies.
Acknowledgements.
@@
The authors thank Mengjie Yu, Avik Dutt, and Sri Krishna Vadlamani for fruitful discussions. This work was supported by funding from the DARPA INSPIRED program.
Appendix A Analytical formula derivation
In this section, we analytically derive the squeezing and SHG dynamics generated in a PPLN device. Let $\hat{a}_{\omega}$ and $\hat{a}_{2\omega}$ denote the field operators at frequencies $\omega$ and $2\omega$, respectively. Both nonlinear interactions arise from the same three–wave mixing Hamiltonian at PPLN position $z$, given by
$\hat{\mathcal{H}}=\frac{\hbar g}{2i}\left(\hat{a}_{\omega}^{2}\,\hat{a}_{2\omega}^{\dagger}\,e^{i\Delta kz}-\hat{a}_{\omega}^{\dagger\,2}\,\hat{a}_{2\omega}\,e^{-i\Delta kz}\right),$ (5)
where $g$ is the nonlinear coupling strength, $\Delta k=2k_{\omega}-k_{2\omega}$ is the wave-vector mismatch.
Thus, the temporal evolution of the operators are governed by the Heisenberg–Langevin equations
$\frac{d\,\hat{a}_{2\omega}}{dz}$ $=-\frac{g}{2}\,\hat{a}^{2}_{\omega}\,e^{i\Delta kz}-\frac{\Gamma_{2\omega}}{2}\,\hat{a}_{2\omega}+\sqrt{\Gamma_{2\omega}}\,\hat{v}_{2\omega}$ (6)
$\frac{d\,\hat{a}_{\omega}}{dz}$ $=g\,\hat{a}_{\omega}^{\dagger}\,\hat{a}_{2\omega}\,e^{-i\Delta kz}-\frac{\Gamma_{\omega}}{2}\,\hat{a}_{\omega}+\sqrt{\Gamma_{\omega}}\,\hat{v}_{\omega},$
where $\Gamma_{2\omega}$ and $\Gamma_{\omega}$ are the PPLN waveguide propagation losses at the corresponding modes, $\hat{a}_{2\omega}$ and $\hat{a}_{\omega}$ are the field operators, $\hat{v}_{2\omega}$ and $\hat{v}_{\omega}$ are the vacuum noise operators. The field and vacuum operators satisfy the commutation relations:
$\left[\hat{a}_{\nu},\hat{a}^{\dagger}_{\mu}\right]=\left[\hat{v}_{\nu},\hat{v}^{\dagger}_{\mu}\right]=\delta_{\nu\mu}\quad,\quad\langle\hat{a}^{\dagger}_{\nu}\hat{a}_{\mu}\rangle=\langle\hat{v}^{\dagger}_{\nu}\hat{v}_{\mu}\rangle=0,$ (7)
for $\nu,\,\mu\in\{2\omega,\omega\}$. Although both squeezing and SHG processes originate from the same interaction Hamiltonian, the relevant approximations and boundary conditions differ. In the following subsections, we solve the Heisenberg–Langevin equation of SHG and squeezing.
A.1 Second-harmonic generation
In the SHG process, the pump field at frequency $\omega$ is treated as a strong, undepleted classical drive such that $\langle\hat{a}_{\omega}\rangle\approx A_{\omega}$ remains constant along the waveguide and the non-linear coupling strength of SHG is
$g_{\text{SHG}}=\frac{2\omega\,d_{\text{eff}}}{n_{2\omega}\,c}\sqrt{\frac{\hbar\omega\,S_{2\omega}}{\epsilon_{0}\,S^{2}_{\omega}\,L}},$
where $d_{\text{eff}}=2d_{33}/\pi$ is the effective nonlinear coefficient under quasi-phase-matching, $d_{33}$ is the nonlinear optical coefficient, $c$ is the speed of light, $\epsilon_{0}$ denotes the vacuum permittivity, $n_{2\omega}$ is the refractive index of LiN at $2\omega$, $S_{2\omega}$ and $S_{\omega}$ are the effective mode areas at the corresponding modes, and $L$ is the poling length of PPLN.
Under this approximation, the second-harmonic mode $\hat{a}_{2\omega}$ obeys a driven linear equation, and its mean-field solution at position $z=L$ is
$\langle\hat{a}_{2\omega}(L)\rangle$ $\approx-\frac{g_{\text{SHG}}\,A_{\omega}^{2}}{2}\,\int_{0}^{L}\,dz^{\prime}\,e^{\left(i\Delta k+\frac{\Gamma_{2\omega}}{2}\right)\,(z^{\prime}-L)}$ (8)
$=-\frac{g_{\text{SHG}}\,A_{\omega}^{2}}{2}\,\left(\frac{1-e^{-\left(i\Delta k+\frac{\Gamma_{2\omega}}{2}\right)\,L}}{i\Delta k+\Gamma_{2\omega}/2}\right).$
where $\Delta k$ is the wave-vector mismatch, and $\Gamma_{2\omega}$ is the loss of the generated second-harmonic field.
$|\langle\hat{a}_{2\omega}(L)\rangle|^{2}=g_{\text{SHG}}^{2}\,A_{\omega}^{4}\,\left(\frac{1+e^{-\Gamma_{2\omega}L}-2e^{-\frac{\Gamma_{2\omega}L}{2}}\cos{\left(\Delta kL\right)}}{4\Delta k^{2}+\Gamma^{2}_{2\omega}}\right).$ (9)
To quantify the overall frequency-conversion performance, we express the generated second-harmonic signal in terms of optical power. For a mode at frequency $\Omega\in\{\omega,2\omega\}$, the optical power is related to its electric-field amplitude $E_{\Omega}$ by
$P_{\Omega}=\frac{n_{\Omega}\,c\,\epsilon_{0}}{2}\,|E_{\Omega}|^{2}S_{\Omega}.$ (10)
The fields $E_{2\omega}$ and $E_{\omega}$ are obtained by normalizing the dimensionless amplitudes $\langle\hat{a}_{2\omega}(L)\rangle$ and $A_{\omega}$ to the corresponding single-photon electric fields. In particular, they are related through
$E_{2\omega}=\langle\hat{a}_{2\omega}(L)\rangle\sqrt{\frac{4\hbar\omega}{n_{2\omega}^{2}\epsilon_{0}\,S_{2\omega}L}}\quad,\;E_{\omega}=A_{\omega}\sqrt{\frac{2\hbar\omega}{n_{\omega}^{2}\epsilon_{0}\,S_{\omega}L}}.$ (11)
In this way, $\left|\langle\hat{a}_{2\omega}(L)\rangle\right|^{2}$ and $\left|A_{\omega}\right|^{2}$ directly represent the mean photon numbers at $2\omega$ and $\omega$, while $E_{2\omega}$ and $E_{\omega}$ carry the proper physical units required in the power expression of Eq. (10).
Using this relation, the on-chip second-harmonic power $P^{(\text{ON})}_{2\omega}$ referenced to the off-chip pump power $P_{\omega}^{(\text{OFF})}$ leads to the on-chip SHG efficiency [17]:
$\mathcal{E}^{(\varepsilon_{2},\,\Delta k)}_{\text{ON}}(x_{2})$ $\equiv\frac{P^{(\text{ON})}_{2\omega}}{\left(P_{\omega}^{(\text{OFF})}\,L\right)^{2}}=\frac{2\,n_{2\omega}\,S_{2\omega}}{n^{2}_{\omega}\,\epsilon_{0}\,c\,L^{2}\,S^{2}_{\omega}}\left|\frac{E_{2\omega}}{E_{\omega}^{(\text{OFF})\,2}}\right|^{2}$ (12)
$=\beta\,x_{2}^{2}\,\varepsilon_{2}^{2}\,\left(\frac{1+e^{-\Gamma_{2\omega}L}-2e^{-\frac{\Gamma_{2\omega}L}{2}}\cos{\left(\Delta kL\right)}}{\Delta k^{2}+\Gamma^{2}_{2\omega}/4}\right),$
where
$\beta=\frac{8\pi^{2}d^{2}_{\text{eff}}}{n_{2\omega}\,n_{\omega}^{2}\,c\,\epsilon_{0}\,L^{2}\,\lambda_{\omega}^{2}}\frac{S_{2\omega}}{S_{\omega}^{2}},$
i.e., $\lambda_{\omega}=2\pi c/\omega$. Here, $E_{\omega}^{(\text{OFF})}$ denotes the off-chip pump electric field, and $n_{\omega}$ is the refractive index of LiN at $\omega$. After coupling into the PPLN waveguide, it becomes $E_{\omega}=x_{2}\,\varepsilon_{2}\,E_{\omega}^{(\text{OFF})}$ with tunable coupling efficiency $x_{2}$ and pump-power fluctuation rate $\varepsilon_{2}$.
A.2 Squeezed light
Squeezed-light generation in the PPLN waveguide arises from the SPDC process in which the $2\omega$ pump is converted into photon pairs at $\omega$. Under strong pumping, the $2\omega$ mode is treated as a classical undepleted drive, $\langle\hat{a}_{2\omega}\rangle\approx A_{2\omega}$, which leads to
$\frac{d\,\hat{a}_{\omega}}{dz}=r\,e^{-i\Delta kz}-\frac{\Gamma_{\omega}}{2}\,\hat{a}_{\omega}+\sqrt{\Gamma_{\omega}}\,\hat{v}_{\omega},$ (13)
where $r\equiv g_{\text{SQZ}}\,A_{2\omega}\,L$ denotes the on-chip squeezing level, coupling strength
$g_{\text{SQZ}}=\frac{\omega\,d_{\text{eff}}}{n_{\omega}\,c}\sqrt{\frac{4\hbar\omega}{n_{2\omega}^{2}\,\epsilon_{0}\,S_{2\omega}\,L}}.$
Unlike SHG, Eq. (13) admits an additional simplification. In a single-pass waveguide, both the nonlinear gain and the propagation loss accumulate gradually, and over an infinitesimal segment $dz$ they act on the same mode in a way that does not meaningfully interfere. To the first order $\mathcal{O}(dz)$, it makes no physical difference whether a small amount of loss occurs before or after a small amount of squeezing. Consequently, the entire distributed evolution can be modeled as an ideal Bogoliubov transformation followed by a single effective loss channel.
Under this approximation, the squeezing unitary operation is described by
$\hat{U}(L)$ $=\exp{\left[\frac{r}{2}\,\left(\hat{a}^{2}_{\omega}\,\left(\int_{0}^{L}\,dz\,e^{i\,\Delta k\,z}\right)-\hat{a}^{\dagger\,2}_{\omega}\,\left(\int_{0}^{L}\,dz\,e^{-i\,\Delta k\,z}\right)\right)\right]}$ (14)
$=\exp{\left[\frac{r}{2}\,\left(\hat{a}^{2}_{\omega}\,\left(\frac{e^{i\,\Delta k\,L}-1}{i\,\Delta\,k}\right)-\hat{a}^{\dagger\,2}_{\omega}\,\left(\frac{e^{-i\,\Delta k\,L}-1}{-i\,\Delta\,k}\right)\right)\right]}$
$=\exp{\left[\frac{r^{\prime}}{2}\,\left(\hat{b}^{2}_{\omega}-\hat{b}^{\dagger\,2}_{\omega}\right)\right]},$
where $\phi=\text{Arg}\left(\frac{e^{i\,\Delta k\,L}-1}{i\,\Delta\,k}\right)$, $r^{\prime}=r\left|\text{sinc}\left(\frac{\Delta k\,L}{2}\right)\right|$ and $\hat{b}_{\omega}=\hat{a}_{\omega}\,e^{i\phi/2}$. Loss is treated as a single amplitude-damping channel with total transmission $\eta(L)=e^{-\Gamma_{\omega}L}$. The output field operator in the rotating frame ($\hat{b}_{\omega}$) is therefore
$\hat{b}_{\omega}(L)$ $\approx\sqrt{\eta(L)}\,\left(\hat{U}^{\dagger}(L)\,\hat{b}_{\omega}\,\hat{U}(L)\right)+\sqrt{1-\eta(L)}\,\hat{v}_{\omega}$ (15)
$=e^{-\frac{\Gamma_{\omega}\,L}{2}}\,\left(\hat{b}_{\omega}\,\cosh{r^{\prime}}-\hat{b}^{\dagger}_{\omega}\,\sinh{r^{\prime}}\right)+\sqrt{1-e^{-\Gamma_{\omega}\,L}}\,\hat{v}_{\omega}.$
Evaluating the field at the output $z=L$ gives
$\hat{q}_{\omega}(L)$ $=e^{-\frac{\Gamma_{\omega}\,L}{2}}\,\left(\hat{b}_{\omega}+\hat{b}^{\dagger}_{\omega}\right)\,e^{-r^{\prime}}+\sqrt{1-e^{-\Gamma_{\omega}\,L}}\,\left(\hat{v}_{\omega}+\hat{v}^{\dagger}_{\omega}\right)$ (16)
$\hat{p}_{\omega}(L)$ $=e^{-\frac{\Gamma_{\omega}\,L}{2}}\,\left(\frac{\hat{b}_{\omega}-\hat{b}^{\dagger}_{\omega}}{i}\right)\,e^{r^{\prime}}+\sqrt{1-e^{-\Gamma_{\omega}\,L}}\,\left(\frac{\hat{v}_{\omega}-\hat{v}^{\dagger}_{\omega}}{i}\right),$
using the conventions $\hat{q}_{\omega}=\hat{a}_{\omega}+\hat{a}_{\omega}^{\dagger}$ and $\hat{p}_{\omega}=(\hat{a}_{\omega}-\hat{a}_{\omega}^{\dagger})/i$. Since the input and vacuum modes have zero mean, $\langle\hat{q}_{\omega}(L)\rangle=\langle\hat{p}_{\omega}(L)\rangle=0$.
The quadrature variances follow directly [46]:
$\Delta q^{2}_{\omega}(L)$ $=e^{-\Gamma_{\omega}\,L}\,e^{-2\,r^{\prime}}+1-e^{-\Gamma_{\omega}\,L}<1\quad\quad\text{(squeezing)}$ (17)
$\Delta p^{2}_{\omega}(L)$ $=e^{-\Gamma_{\omega}\,L}\,e^{2\,r^{\prime}}+1-e^{-\Gamma_{\omega}\,L}>1\quad\text{(anti-squeezing)}.$
To relate the squeezing level of the field to the pump power, we express the pump amplitude as
$A_{2\omega}$ $=\sqrt{\frac{2\,x_{1}\,\varepsilon_{1}\,P^{(0)}_{2\omega}}{n_{2\omega}\,c\,\epsilon_{0}\,S_{2\omega}}\left(\frac{n_{2\omega}^{2}\,\epsilon_{0}\,S_{2\omega}\,L}{4\,\hbar\,\omega}\right)}=\sqrt{\frac{n_{2\omega}\,L\,x_{1}\,\varepsilon_{1}\,P^{(0)}_{2\omega}}{2\,\hbar\,\omega\,c}},$ (18)
where $P^{(0)}_{2\omega}$ is the amplitude of the off-chip pump power, $x_{1}$ is the tunable off-chip coupling efficiency, and $\varepsilon_{1}$ is the pump fluctuation rate. As a result, the resulting on-chip squeezing level is
$\mathcal{S}^{(\varepsilon_{1},\,\Delta k)}_{\text{ON}}$ $(x_{1})=$ (19)
$-\log\left(e^{-\Gamma_{\omega}L}e^{-2\sqrt{x_{1}\varepsilon_{1}\alpha P^{(0)}_{2\omega}}\,\left|\text{sinc}\left(\frac{\Delta kL}{2}\right)\right|}+1-e^{-\Gamma_{\omega}L}\right),$
where $\alpha=\sqrt{n_{2\omega}\,L/2\hbar\omega c}$.
Appendix B Efficiency thresholds for off-chip squeezing
@@
On-chip squeezing levels approaching $15\penalty 10000\ \text{dB}$ are particularly relevant for applications such as CV quantum error correction and precision quantum metrology [26, 39]. As shown in Fig. 5(a), achieving this on-chip benchmark requires an input coupling efficiency $\eta^{(2\omega)}_{1}\gtrsim 0.65$, which is attainable using current photonic integration technologies. For example, edge-coupled LiN modulators have demonstrated per-interface coupling efficiencies below $0.5\penalty 10000\ \text{dB}$ (corresponding to efficiencies $90\penalty 10000\ \%$) at $1550\penalty 10000\ \text{nm}$ [52], indicating that such coupling efficiencies are experimentally feasible.
@@
Simulation parameters $d_{33}$ $19.5$ pm/V Thin-film LiN [17] nonlinear coefficient $L$ 11 mm  Poling length of PPLN $\omega$ $193$ THz Pumping frequency (or $2\omega$)  (or $386$ THz)   at $1550$ nm (or $775$ nm) $n_{\omega}$ $2.14$  Refractive index of LiN (or $n_{2\omega}$) (or $2.2$)  at $\omega$ (or $2\omega$) $\mathcal{S}_{\omega}$ 11 µm2 Mode area of the field (or $\mathcal{S}_{2\omega}$)  (or $0.8$ µm2) in PPLN waveguide $\Gamma_{\omega}$ 55 dB/m [53] Propagation loss (or $\Gamma_{2\omega}$) (or $10$ dB/m) $\sigma_{\text{sqz}}$ $0.1\penalty 10000\ \text{dB}$  Squeezing (or SHG) (or $\sigma_{\text{shg}}$)    (or $20\penalty 10000\ \text{\%/W/cm${}^{2}$}$)   measurement noise STD Off-chip average $P_{2\omega}$ $100$ mW pump power in the squeezing process Half range of power $\delta_{1}$ $0.5\penalty 10000\ \%$ fluctuation rate in (or $\delta_{2}$) squeezing (or SHG) $\sigma_{\Delta k}$ 0.36 m-1 Wave-vector mismatch STD Linear transmission $\sigma_{\text{C}}$ $7\penalty 10000\ \%$ measurement STD with conventional approach
Table 1: Monte Carlo simulation parameters of the PPLN waveguide in squeezing and SHG processes. STD: standard deviation.